# The Bayesian Separation Principle for Data-driven Control

### v1

@@ VII Conclusion
In this paper we have explored the existence and role of a separation principle for control and identification. By focusing on model predictive control, we have clarified that the separation principle holds universally in a Bayesian setting, i.e. regardless of data size, and asymptotically in Fisherian setting. We further discussed how to cast MPC in a non-parametric Gaussian regression framework and how to exploit the Bayesian separation principle to establish explicit, computationally tractable expressions for the control cost and optimal input sequence. The controllers designed by leveraging this approach exhibit superior performance compared to traditional design techniques, as illustrated by our numerical experiments.
Appendix
Notes on the proof of Proposition 1
@@
In the setting of dynamic systems, we assume that the model (24) for $f$ is not influenced by the past input-output data contained in $z^{-}$. For this reason, in the next equations conditioning on $z^{-}$ is omitted and the Gaussian prior describes all the system randomness before seeing only the data in $\mathcal{D}$. Since $u$ is a fixed deterministic input, $Y$ can be also used in place of $\mathcal{D}$.
@@
$C=\left[\begin{array}[]{cc}\mathbb{K}&\Gamma\\ \Gamma^{\top}&K(z^{*},z^{*})\end{array}\right].$
@@
As for the the marginal probability density function of $Y$ reported in (30), it can be easily obtained using the Laplace integral to integrate out $\bar{f}^{*}$ from $p(\bar{f}^{*},Y)$ [39]. But an even simpler argument is to note that the minus log of the joint density reported in (Notes on the proof of Proposition 1) is the same as if we had started from the model measurements $Y=M\bar{f}^{*}+E$ where $\bar{f}^{*}$ is the Gaussian random field $f$ sampled on deterministic input locations that do not depend on $Y$ itself. This interpretation leads to a problem equivalent to the original one, just for the purposes of marginalization w.r.t $\bar{f}^{*}$, where the vector $Y$ is Gaussian with covariance $\Sigma=K+\gamma I$ [^5] This is certainly not true starting from our original model where $M$ also depends on $Y$, but again we emphasise that this is irrelevant for our integration problem.. This immediately gives (Notes on the proof of Proposition 1) as the result of the integral.
@@
Notes on the proof of Proposition 2
Proof of Proposition 2 is just a specialization of that obtained in [22] where the more complex case of time-varying input locations probability distributions is treated. The result in [22] however holds only if the kernel under study admits (at least) one expansion of the type
$K(x,a)=\sum_{k}\lambda_{k}h_{k}(x)h_{k}(a),\quad\zeta_{k}\geq 0$
with
$\sum_{k}\lambda_{k}<\infty$ (41)
and
$\sup_{x\in X}|h_{k}(x)|<M<+\infty$ (42)
with $M$ independent of $k$. Over any compact domain, such condition is clearly fulfilled by the spline kernel (17), including that of second-order, and by any kernel generated by a number finite of basis functions, like the polynomial (21). It now suffices showing that this kind of expansion exists also for the Gaussian kernel to be certain of the correctness of Proposition 2 here reported.
Let us now consider the more general version of the Gaussian kernel in (20) which depends on the diagonal matrix $D$. It consists of the product of basic Gaussian kernels having different kernel widths $\eta$. Its basis functions and expansion coefficients are thus the (tensor) product of the $\phi_{k}$ and $\zeta_{k}$ reported in (15) and (18). So, one has just to focus on the Gaussian kernel over $\mathbb{R}^{2}$ to check if (41) and (42) hold.
Using [32, Eq. 15] one obtains that the Gaussian kernel has the following expansion over any subset $X$ of $\mathbb{R}^{2}$:
$\exp\big{(}-\frac{(x-t)^{2}}{\eta}\big{)}$
$=\exp\big{(}-\frac{x^{2}}{\eta}\big{)}\exp\big{(}-\frac{t^{2}}{\eta}\big{)}\sum_{k=0}^{+\infty}\frac{(2/\eta)^{k}}{k!}x^{k}t^{k}.$
Without loss of generality let $X={\mathcal{Q}}_{N}$, where ${\mathcal{Q}}_{N}$ is the square of sides $[0,N]$ which contains any compact set $X\subset{\mathbb{R}}^{2}$ for $N$ large enough. Let also $\eta=1$, otherwise we can replace $K(x,t)$ by $K(ax,at)$, with $a>0$, without modifying the properties obtained below. It holds that
$K(x,t)=\sum_{k=0}^{+\infty}\ \psi_{k}(x)\psi_{k}(t)$
with
$\psi_{k}(x):=\sqrt{\frac{2^{k}}{k!}}x^{k}e^{-x^{2}}.$
The non-negative function $\psi_{k}(x)$ has a local maximum at $x=\sqrt{\frac{k}{2}}$ larger than $N$ if $k$ is large enough. So for large $k$ its maximum value is attained at the boundary $x=N$ and turns out
$\sqrt{\frac{2^{k}}{k!}}N^{k}e^{-N^{2}}.$
Let us now consider $h_{k}(x):=k\psi_{k}(x)$, so that
$K(x,t)$ $=$ $\sum_{k=0}^{+\infty}\ \frac{1}{k^{2}}h_{k}(x)h_{k}(t)$
::=:= $\sum_{k=0}^{+\infty}\ \lambda_{k}h_{k}(x)h_{k}(t).$
Clearly, one has $\sum_{k=0}^{+\infty}\lambda_{k}<+\infty$ and the maximum of $h_{k}(x)$, for $k$ large enough, is given by
$k\sqrt{\frac{2^{k}}{k!}}N^{k}e^{-N^{2}}.$
It suffices to show that
$\lim_{k\rightarrow+\infty}\ k\sqrt{\frac{2^{k}}{k!}}N^{k}e^{-N^{2}}=0$
to prove that $h_{k}(x)$’s are uniformly bounded. By Stirling’s formula, for $k\geq k_{0}$ one has
$k\sqrt{\frac{2^{k}}{k!}}N^{k}e^{-N^{2}}\simeq\frac{e^{-N^{2}}k}{\sqrt[4]{2\pi k}}\left(\frac{2eN^{2}}{k}\right)^{\frac{k}{2}}$
$=\frac{e^{-N^{2}}k}{\sqrt[4]{2\pi k}}\ e^{-k\left(\frac{ln(k)}{2}-\frac{ln(2eN^{2})}{2}\right)}$
$\leq\frac{e^{-N^{2}}k}{\sqrt[4]{2\pi k}}\ e^{-k\left(\frac{ln(k_{0})}{2}-\frac{ln(2eN^{2})}{2}\right)}.$
Then, choosing $k_{0}>2eN^{2}$, there exist $A,B>0$ leading to the upper bound
$Ak^{\frac{3}{4}}e^{-Bk}$
which tends to zero as $k$ grows to infinity.
It only remains to prove that convergence holds over a set of functions which can approximate arbitrarily well any continuous function if spline or Gaussian kernels are used. For this purpose, recall that the kernel operator is
[…]
and defines also a set of eigenfunctions $\phi_{i}$ satisfying
$L_{p}[\phi_{i}]=\lambda_{i}\phi_{i}$
which span all the RKHS. According to Assumption 1 in [22], convergence holds for any system $f$ in $\mathcal{H}$ which satisfies the smoothness condition
$\|L_{p}^{-r}f\|_{p}<\infty,\quad\frac{1}{2}<r\leq 1.$ (44)
From [14, Theorem 3] and subsequent discussion, it is easy to see that such subset corresponds exactly to the RKHS still generated by the eigenfunctions $\phi_{i}$. Precisely, any function $f$ admits the representation
$f(x)=\sum_{i}c_{i}\phi_{i}(x)$
with the expansion coefficients that meet the constraint
$\sum_{i}\frac{c_{i}^{2}}{\lambda_{i}^{2r}}.$
Since spline and Gaussian kernels are universal, the $\phi_{i}$ form also a basis of the space of continuous functions and this completes the proof.

### v2

@@ 8 Conclusion
In this paper, we explored the existence and role of a separation principle between identification and control design. By focusing on model predictive control, we argued that such separation principle holds asymptotically in the number of data in a Fisherian setting and also for finite data in a Bayesian setting. We further discussed how to cast MPC within a non-parametric Gaussian regression framework and how to exploit the separation principle to establish explicit, computationally tractable expressions for the control cost and optimal input sequence. The control inputs designed by leveraging this approach may exhibit superior performance compared to both nominal, uncertainty-agnostic techniques and alternative uncertainty-aware formulations, as illustrated by our numerical experiments. Throughout the paper we also provided precise definitions for commonly used terminology in the data-driven control literature.
Overall, this work thus points out important advantages related to the use of the Bayesian setting for control design. One could however argue that the price to pay is the need to introduce and believe in the prior. This aspect, however, is greatly mitigated by the introduction of very general priors typical of Gaussian regression which include just regularity information on the input-output map. In addition, the hyperparameters that regulate the smoothness level are tuned from the data, providing a form of robustness. There are also routes to further enhance such robustness, e.g. resorting to the even more sophysticated frequentist Bayesian framework [5, 2] which may account for misspecified priors when calculating uncertainty bounds. Possible developments of this work may thus include a comparison between different statistical paradigms for the design of controllers including Bayes, Fisher and frequentist Bayes.
Appendix: Proof of Proposition 1
@@
The past input-output data preceding in time those contained in $\mathcal{D}$ are contained in the vector denoted by $z^{-}$ and are used to define the initial conditions of our dynamic models. We assume that the model (23) for $f$ is not influenced by these past input-output data. Therefore, in the following conditioning on $z^{-}$ is omitted and the Gaussian prior on $f$ describes all system randomness before seeing only the data in $\mathcal{D}$. Since $u$ is a fixed deterministic input, we can also leave its dependence implicit and $Y$ can be used instead of $\mathcal{D}$ to define the $\sigma$-algebra on which $f$ is projected.
@@
$C=\lambda\left[\begin{array}[]{cc}\mathbb{K}&\Gamma\\ \Gamma^{\top}&K(z^{*},z^{*})\end{array}\right].$
@@
As for the the marginal probability density function of $Y$ reported in (29), it can be easily obtained using the Laplace integral to integrate out $\bar{f}^{*}$ from $p(\bar{f}^{*},Y)$ [36]. But an even simpler argument is to note that the minus log of the joint density reported in (Appendix: Proof of Proposition 1) is the same as if we had started from the model measurements $Y=M\bar{f}^{*}+E$ where $\bar{f}^{*}$ is the Gaussian random field $f$ sampled on deterministic input locations that do not depend on $Y$ itself. This interpretation leads to a problem equivalent to the original one, just for the purposes of marginalization w.r.t $\bar{f}^{*}$, where the vector $Y$ is zero-mean Gaussian [^5] This is certainly not true starting from our original model where $M$ also depends on $Y$, but again we emphasise that this is irrelevant for our integration problem. with covariance $\Sigma_{y}=\lambda K+\sigma^{2}I$. This immediately gives (Appendix: Proof of Proposition 1) as the result of the integral.